OpenStudy (arihangdu):

Fe -59 has a half life of 45.1 days. How old is an iron nail if the Fe-59 content is 25% that of a new sample of iron? Please show all calculation.

OpenStudy (arihangdu):

Will medal and fan as soon as possible.

OpenStudy (cuanchi):

One half life is the time that takes to decrease the amount of the nucleoide in half of the original amount. That means that in 45.1 days (one half life) you will have 50% of the original amount of Fe-59 compared with the new sample of iron In two half life (45.1 days x 2) you will have 50% of the 50%, that means 25% of the new sample contents. What do you think is the answer?

OpenStudy (arihangdu):

25%=25/100 =1/4. (1/2)n=1/4.so, n=2 half lives. 45.1 days x 2 half lives=90.2 days . Right?

OpenStudy (arihangdu):

Is this correct? My friends. I'm not sure.

OpenStudy (arihangdu):

Please help me to check my answer

OpenStudy (cuanchi):

Yes that is Perfect of a Master!!!!

OpenStudy (arihangdu):

Hank you my friend.